The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Bryan C. Dunne Co-Editors Eva K. Grebel and You-Hua Chu

                                 mcnews@astro.uiuc.edu
                 http://www.astro.uiuc.edu/projects/mcnews/MCNews.html
                         http://www.astro.uni-bonn.de/~mcnews/

No. 65 July 30, 2002

Contents

Letter From The Editor  1
Abstracts of 2 Refereed Papers  2
Abstracts of 3 Non-Refereed Papers 3
Meeting Announcement  5

Letter From The Editor

Dear Colleagues,

The Magellanic Clouds Working Group is proud to announce the availability ot the complete MC News archive in L^{A}TEX, Postscript, PDF, and HTML formats. The entire archive of issues is available for

download from our Website at http://www.astro.uiuc.edu/projects/mcnews/MCNews.html. Look

for more enhancements to our Website in the future!

Best Wishes, Bryan Dunne

Abstracts of Refereed Papers

Blue Variable Stars from the MACHO database I: Photometry and Spectroscopy of the LMC sample

S. C. Keller^{1}, M. S. Bessell^{2}, K. H. Cook^{1} and M. Geha^{3}

1 IGPP/LLNL, L-413, PO Box 808, Livermore CA 94550, USA 2 RSAA, Mt. Stromlo Obs., Weston Creek, ACT 2611, Australia 3Dept. of Astronomy and Astrophysics, UCO/Lick Observatory, Univ. of California, Santa Cruz, CA 95064, USA

We present the photometric properties of 1279 blue variable stars within the LMC. Photometry is derived from the MACHO database. The lightcurves of the sample exhibit a variety of quasi-periodic and aperiodic outburst behavior. A characteristic feature of the photometric variation is that the objects are reddest when at maximum outburst. A subset of 102 objects were examined spectroscopically. Within this subset, 91% exhibited Balmer emission in at least one epoch, in some cases with spectacular spectral variability. The variability observed in the sample is consistent with the establishment and maintenance of the Be phenomenon.

Accepted by: Astronomical Journal

For preprints, contact skeller@igpp.ucllnl.org

Also available from the URL astro-ph/0206444

The Low End of the Initial Mass Function in Young Clusters: II. Evidence for Primordial Mass Segregation in NGC 330 in the SMC

M. Sirianni^{1}, A. Nota^{2}, G. De Marchi^{2}, C. Leitherer^{2} and M. Clampin^{2}

1 The Johns Hopkins University, Baltimore MD 21218, USA 2 Space Telescope Science Institute, Baltimore MD 21218, USA

As part of a larger program aimed at investigating the universality of the initial mass function (IMF) at low masses in a number of young clusters in the LMC/SMC, we present a new study of the low end of the stellar IMF of NGC 330, the richest young star cluster in the SMC, from deep broadband V and I images obtained with HST/WFPC2. We detect stars down to a limiting magnitude of m_{555} = 24.9 which corresponds to stellar masses of ∼ 0.8 M_{⊙} at the distance of the SMC. A comparison of the cluster color-magnitude diagram with theoretical evolutionary tracks indicates an age of ∼ 30 Myr for NGC 330, in agreement with previous published results. We derive the cluster luminosity function, which we correct for background contamination using an adjacent SMC field, and construct the mass function (MF) in the 1−7 M_{⊙} mass range. Given the young cluster age, the MF can well approximate the IMF. We find that the IMF in the central cluster regions (within 30”) is well reproduced by a power-law with a slope consistent with Salpeter’s. In addition, the richness of the cluster allows us to investigate the IMF as a function of radial distance from the center. We find that the IMF becomes steeper at increasing distances from the cluster center (between 30” and 90” ), with the number of massive stars (> 5 M_{⊙}) decreasing from the core to the outskirts of the cluster five times more rapidly than the less massive objects (≃ 1 M_{⊙}). We believe the observed mass segregation to be of primordial nature, rather then dynamical, since the age of NGC 330 is ten times shorter than the expected relaxation time of the cluster.

Accepted by: Astrophysical Journal

For preprints, contact sirianni@pha.jhu.edu
Also available from the URL http://www.journals.uchicago.edu/ApJ/future.html

Abstracts of Non-Refereed Papers

Three aspects of red giant studies in the Magellanic Clouds

M.-R.L. Cioni

ESO, Karl–Schwarzschild–str. 2, 85748 Garching bei Muenchen, Germany

There are three important aspects concerning the study of the red giant and in particular of the asymptotic giant branch (AGB) stars in the Magellanic Clouds. These are: the surface distribution, the luminosity function and the variability. The spatial distribution of AGB stars is an efficient tool to study the structure of the galaxies and their metalicity by analysing the ratio between carbon– and oxygen–rich AGB stars. The shape of the luminosity function carries informations about the star formation rate in the Clouds and it can be mathematically related to their history. Most AGB stars vary their magnitude in a few to several hundred years time; the one epoch DENIS magnitudes for both Large and Small Magellanic Cloud AGB stars outline the same relations as a function of period.

To appear in: Mass-Losing Pulsating Stars and their Circumstellar Matter , Y. Nakada & M. Honma (eds) Kluwer ASSL series For preprints, contact mcioni@eso.org

Observational Constraints on Massive Star Evolution

Phil Massey^{1}

1 Lowell Observatory

In this paper, I discuss the observational quantities that are useful for judging the successes and failures of current massive star evolutionary theory. The galaxies of the Local Group can serve as important laboratories for providing these diagnostics, as their metallicities vary by a factor of 10. We find that the evolutionary tracks do a good job of matching the distribution of stars in the H-R diagram during the main-sequence phase. However, none of the models produce RSGs that are as cool and as bright as what is observed. The relative number of WC and WN stars is a strong function of metallicity, and the Padova and Geneva “normal mass-loss” models do a reasonably good job of matching the observations at low metallicities, but predict too few WCs at higher metallicity. The “enhanced” mass-loss models of the Geneva group do not match the observations at all. New data is providing excellent statistics on the number of RSGs in these nearby galaxies, and the number ratio of RSGs to WRs is also an extremely sensitive function of metallicity. None of the models reproduce the trend of the RSG/WR ratio with metallicity.

To appear in: IAU212

Available from the URL ftp://ftp.lowell.edu/pub/massey/iau212massey.ps.gz

Primordial or Dynamical Mass Segregation in Young LMC Clusters?

Richard de Grijs^{1}, Gerry Gilmore^{1}, and Rachel Johnson^{1,2}

1 Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge CB3 0HA, UK 2 European Southern Observatory, Casilla 19001, Santiago 19, Chile

We present the detailed analysis of Hubble Space Telescope observations of the spatial distributions of different stellar species in two young compact star clusters in the Large Magellanic Cloud, NGC 1805 and NGC 1818. Based on a comparison of the characteristic relaxation times in their cores and at their half-mass radii with the observed degree of mass segregation, it is most likely that significant primordial mass segregation was present in both clusters, particularly in NGC 1805. Both clusters were likely formed with very similar initial mass functions.

To appear in: New Horizons in Globular Cluster Astronomy, June 2002, Padova (Italy), eds. Piotto G., Meylan G., Djorgovski S.G., Riello M., A.S.P.: San Francisco For preprints, contact grijs@ast.cam.ac.uk

Also available from the URL http://arXiv.org/abs/astro-ph/?0207247

Meeting Announcements

The Outer Edges of Dwarf Irregular Galaxies: Stars and Gas

Venue: Flagstaff, Arizona, USA Dates: October 10 to 11, 2002 Contact: Deidre Hunter and Sally Oey Address: Lowell Observatory, 1400 W. Mars Hill Road, Flagstaff AZ 86001, USA Phone: +1-928-774-3358

E-mail: lowell02@lowell.edu
URL: http://www.lowell.edu/Workshops/Lowell02/

From the meeting abstract: We have in recent years come to view dwarf galaxy evolution in the broader context of the cosmic evolution of large-scale structure. Dwarf galaxies, as the putative building blocks of hierarchical galaxy formation, and also as the most numerous galaxies in the Universe, play a central role in cosmic evolution. In particular, the interplay of galactic and intergalactic material around dwarf irregulars must be more extensive than in more massive disk galaxies because of their lower gravitational potential and lower interstellar pressures. The outer regions of dwarf irregular galaxies therefore yield vital clues to the dominant processes in this interaction zone.